What is the speed of a truck that ravels 10 km in 10 minutes?
Answer by texttutoring(324)

D=VT
D is distance, V is velocity (speed), T is time.
Let's work in hours: 10 minutes = 1/6 hours
D=VT
V=D/T
V= 10/(1/6)
V=60km/hour